OpenStudy (anonymous):

If f(x) = −3x + 1, find f(−2).

OpenStudy (anonymous):

\[f(x)=-3x+1\] \[f(-2)=?\] Everytime you see the letter x, replace it with a -2.

OpenStudy (anonymous):

That's what they want you to do. ANd then you would get your answer.

OpenStudy (anonymous):

whats the answer then?

OpenStudy (anonymous):

The term f(x) means a function is written in terms of x. For f(-2), it's the exact same thing except it is written in terms of -2. So, with that mind, whenever you come across a "x", just substitute it with a -2. So, f(-2) = -3 (-2) +1 f(-2) = 6 + 1 f(-2) = 7

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

@swilson, it's not their benefit, They did not learn anything from you said but just to get eha answer. Please do not spill the beans

OpenStudy (anonymous):

@swilson9

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

@username1212 Do you actually understand how to solve these questions? Because if you do, then would you be able to solve this. \[f(x)=7x+1\] \[f(-2)=?\]

OpenStudy (anonymous):

whats the answer for this question ? If f(x) = x + 2, find the value of x if f(x) = 12

OpenStudy (anonymous):

@swilson9 Do you see what I mean...He' just wants the answer.

OpenStudy (anonymous):

@username1212 all you have to do is but 12 in every spot there is an x in the equation. Do you get it now

OpenStudy (anonymous):

@username1212 Please do that on your own using @swilson9 method/approach to your previous question.

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

Look @Azteck he did not get your explanation so i gave him an easier way to understand it. Some people learn form seeing the steps i provided so plz do not @ me any more

OpenStudy (anonymous):

Yeah and now he wants the answer from you again with a question that's exactly similar to the previous one. Good job.

OpenStudy (anonymous):

Read the Code of Conduct and you will see you cannot give answers. You can only guide them to the answer. This isn't yahoo answers.